OpenStudy (anonymous):

Given F(2)=1, F'(2)=9, F(4)=3, F'(4)=7 and G(4)=2, G'(4)=6, G(3)=4, and G'(3)=9, find.... Part 1: H(4) if H(x)=F(G(x)) Part 2: H ' (4) if H(x)=F(G(x)) Part 3: H(4) if H(x)=G(F(x)) Part 4: H ' (4) if H(x)=G(F(x)) Part 5: H ' (4) if H(x)= F(x) ------- G(x) please explain!! :) thank you! :D

terenzreignz (terenzreignz):

Wasn't going to log in, but then I saw you online :> I don't really see why those derivatives are necessary, but if H(x) = F(G(x)) Then shouldn't H(4) simply be F(G(4))? LOL

hartnn (hartnn):

because thats the 1st part of a 5 part question derivatives will be useful in other parts

terenzreignz (terenzreignz):

Yeah, didn't see that part until the last minute lol my bad. But what say you, food lover? :3

OpenStudy (nincompoop):

post all of them at once

OpenStudy (anonymous):

hahaa thanks for logging on :) and yeah what @hartnn said :P ohhh so would i be solving for F(2) since G(4)=2?

terenzreignz (terenzreignz):

Good so far, the answer is then....?

OpenStudy (anonymous):

ermm.. F(2)=1 so H(4)=1 ? and i've edited it with all five parts:)

hartnn (hartnn):

thats correct, H(4) is indeed 1 for 2nd part, you'll need chain rue, know what that is ?

OpenStudy (anonymous):

awesome!! :) it's where you work from the outer function first right? but i'm not too good at it haha :(

OpenStudy (nincompoop):

chain rue is a combination of different rues LUL

OpenStudy (nincompoop):

have you come up with a technique how to approach this problem?

OpenStudy (anonymous):

not really sure how to start it off :(

terenzreignz (terenzreignz):

Chain rule : \[\Large \frac{d}{dx}f\left(g(x)\right)=f'\left(g(x)\right) g'(x)\]

hartnn (hartnn):

will tell you how to do part 2, you try part 3,4 by yourself. chain rule written with other notation: \([F(G(x))]' = F'(G(x)) \times G'(x)\) i think you can find the value of terms on right side, try it

OpenStudy (anonymous):

ohhh okay yeah, I'm not too good with the chain rule... yet :P so F'(G(4)) x G'(4) = F'(2) x 6 = 9 x 6 = 54 ?

OpenStudy (anonymous):

okay @hartnn :) so would part 3 be this? H(4) if H(x)=G(F(x)) H(4)=G(F(4)) =G(3) =4 So for part 3, H(4)=4 ? does that look right?

terenzreignz (terenzreignz):

Ohh you.

OpenStudy (anonymous):

hahah :P

OpenStudy (anonymous):

wait so did I not do part 2 right? :/

hartnn (hartnn):

yes, both 2nd , 3rd part are correct :)

OpenStudy (anonymous):

oh yay!! :) lemme try part 4.. just a moment :)

terenzreignz (terenzreignz):

If I knew hart was going to butt in, I wouldn't have logged on XD LOL JK @hartnn ^_^ Signing off now, anyway, good luck, food lover, and TJ out :)

OpenStudy (anonymous):

lawlzz okay, thanks @terenzreignz :) see you later!!

hartnn (hartnn):

did i just scare him away :O :P :(

OpenStudy (nincompoop):

someone should do a tutorial lesson on calculus

OpenStudy (anonymous):

so would i start part 4 like this? H ' (4) if H(x)=G(F(x)) [G(F(x))]' = G'F(x) * F ' (x) ? and @hartnn i don't think so hahaha :P and @nincompoop, yes! though i would definitely not nominate myself for that job... for obvious reasons lol

hartnn (hartnn):

yep, G'F(x) * F ' (x) G'(F(x)) * F ' (x) , careful with brackets, else you might get confused later or could get a wrong answer..

OpenStudy (anonymous):

ohh okay will keep that in mind :) so now it looks like this? G'(F(4)) * F ' (4) = G'(3) * 7 =9 * 7 =63 so H ' (4) for part 4 equals 63?

hartnn (hartnn):

sure it is :)

OpenStudy (anonymous):

yay!! :) so last part now!! how would this one work? :/

hartnn (hartnn):

know quotient rule ?

OpenStudy (anonymous):

\[\frac{ u'v-uv' }{ v^2 }\] ?

hartnn (hartnn):

yes :) use it

OpenStudy (anonymous):

okay:) so would f(x) = u and g(x)=v ?

hartnn (hartnn):

go on, i will stop u , if u make an error...

OpenStudy (anonymous):

okay:) and then (f ' (4) * g(4)) - (f(4) * (g ' (4)) ------------------------- g(4)^2 ?

OpenStudy (anonymous):

and then (7 * 2) - (3 * 6) ------------- 2^2 = 14 - 18 --------- 4 = -4/4 = -1? so H'(4) for H(x)= f(x)/g(x) for part 5 equals -1? did i do that right?

hartnn (hartnn):

seems correct to me :)

OpenStudy (anonymous):

ahh yay!! and we're done!! thanks so much!!!! :D

hartnn (hartnn):

welcome ^_^